OpenStudy (anonymous):

Angular and linear speed…. A cyclist is riding a bicycle whose wheels have a radius of 10 inches. Suppose he is traveling at 22 miles per hour. (A useful fact: =1mi5280ft .) (a) Find the angular speed of the wheels in radians per minute. (b) Find the number of revolutions the wheels make per minute. Do not round any intermediate computations, and round your answer to the nearest whole number.

jimthompson5910 (jim_thompson5910):

what do you have so far?

OpenStudy (anonymous):

nothing yet… I'm not to sure how to start.

jimthompson5910 (jim_thompson5910):

so the radius is 10 inches? and the speed is 22 mph? or are any values given as fractions or something?

OpenStudy (anonymous):

yes and yes… and no

jimthompson5910 (jim_thompson5910):

Given the radius r = 10, what is the circumference of the bike wheel?

OpenStudy (anonymous):

idk

jimthompson5910 (jim_thompson5910):

you would use the formula C = 2*pi*r C is the circumference pi is approximately 3.14 r is the radius

OpenStudy (anonymous):

so 20pi

jimthompson5910 (jim_thompson5910):

very good

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

the units for the circumference is in inches

OpenStudy (anonymous):

so now what?

jimthompson5910 (jim_thompson5910):

the speed is 22 miles per hour we have inches for the wheel but miles for the speed, so we first have to convert from miles to inches ( 22 miles ) * ( 5280 ft/1 mile) = 22*5280 = 116,160 feet (116,160 feet) * (12 inches/1 ft) = 116,160*12 = 1,393,920 inches ------------------------------------------------------- So we now know that 22 miles = 1,393,920 inches making sense so far?

OpenStudy (anonymous):

yes but my computer in gonna die and i don't have my charger.

jimthompson5910 (jim_thompson5910):

how much time do you have?

jimthompson5910 (jim_thompson5910):

if not a lot, then I can come back later

OpenStudy (anonymous):

its at 2%

OpenStudy (anonymous):

its is making sense though

jimthompson5910 (jim_thompson5910):

anyways, 22 mph = 1,393,920 inches per hour we need to go from inches to hour to inches per min so we multiply by (1 hr/60 min) (1,393,920 inches/1 hr)*(1 hr/60 min) = 1,393,920/60 = 23,232 inches per min

jimthompson5910 (jim_thompson5910):

it's a lot of steps, but we now know 22 mph = 23,232 inches per min

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

One full revolution corresponds to the circumference C so we can say 1 rev = 1 C which is the same as 1 rev = 20pi inches we now take our last result (23,232 inches per min) and multiply it by (1 rev/20pi inches) to get... (23,232 inches/1 min)*(1 rev/20pi inches) = 23,232/20 = 1,161.6 this means that the angular speed is 1,161.6 revolutions per min convert (1,161.6 rev/1 min) to rad/min (1,161.6 rev/1 min)*(2pi rad/1 rev) = 1,161.6*2pi = 2,323.2pi rad/min

jimthompson5910 (jim_thompson5910):

so the angular speed is 2,323.2 radians per min

OpenStudy (anonymous):

ok i understand that…

jimthompson5910 (jim_thompson5910):

sorry I had to log off, but I'm back

jimthompson5910 (jim_thompson5910):

were you able to answer the second part?

OpenStudy (anonymous):

thats fine no worries no i wasn't able to. :(

jimthompson5910 (jim_thompson5910):

Notice how in my steps I had "1,161.6 rev/1 min". It's towards the end. so it makes 1,161.6 revolutions in 1 minute